OpenStudy (anonymous):

A+B=35 B+C=37 A+C=28 Solve for A, B, and C

OpenStudy (anonymous):

A=35-B C=37-C A+C=28=>35-B+37-B=28=>72-2B=28=>2B=44=>B=22 A=13 C=15

OpenStudy (anonymous):

B + C = 37 A + C = 28 ->-A +B = 9 A + B = 35 => 2B = 44 Thus B = 22, A = 35 -22 = 13 C = 28 - 13 = 15